# Unbounded: A Generative Infinite Game of Character Life Simulation

### v1

@@ Abstract
Figure 1: An example of Unbounded. We follow the life of Archibus, the user’s custom wizard character. The user can interact with the generative game using natural language, and Archibus’ hunger, energy and fun meters update accordingly. A spontaneous and unconstrained story unfolds while the user playing, and the character can explore new environments with a myriad of possible actions and unexpected interactions. The game runs in interactive speeds, refreshing every second.
@@ 1 Introduction
In his work “Finite and Infinite Games: A Vision of Life as Play and Possibility” Carse (1986), James P. Carse introduces a distinction between two types of games. Carse defines finite games as those “played for the purpose of winning,” with boundaries, fixed rules, and a definitive endpoint. In contrast, infinite games are “played for the purpose of continuing the play,” with no fixed boundaries and evolving rules. Traditional video games are, inherently, finite games due to the limitations of computer programming and computer graphics. For example, all game mechanics have to be fully pre-defined in the programming language and all graphics assets have to be pre-designed (modulo procedural generation which still grapples with structural limitations). This allows for only a finite, and sometimes predefined, set of actions and paths that can be taken. They also feature predefined rules, boundaries, and win conditions.
@@
Recent advances in generative models have been impressive. We hypothesize that these developments have finally opened up the possibility of creating the first generative infinite video game. Two advancements have made this achievable: (1) large language models (LLMs) that can encode persistent video game mechanics (e.g. interactions with the game environment or characters, character state tracking, object permanence), generate interactive stories, and produce spontaneous, and sometimes emergent, behaviors; and (2) visual generative models capable of producing high-quality images that follow prompts. In this work, we present Unbounded, what we believe to be the first interactive generative infinite game, where game behaviors and outputs are generated by AI models, transcending the constraints of hard-coded systems. Unbounded draws inspiration from sandbox life simulations and digital pet games such as Little Computer People, The Sims and Tamagotchi. It incorporates elements from tabletop roleplaying games like Dungeons & Dragons, which offer unconstrained storytelling experiences that have been unattainable in video games.
@@ 2 Related Work
Controllable text-to-image generation becomes a key research direction in diffusion model applications, enabling diverse ways to guide the generation process. For instance, ControlNet (Zhang et al., 2023) introduces conditioning mechanisms that utilize control signals such as depth maps, poses, edges, and segmentation maps to guide image generation. Other works focus on layout control using bounding boxes to control object placement within the generated images (Li et al., 2023; Shin et al., 2022). Beyond these control signals, another major area of research involves personalization, where the goal is to generate consistent characters (Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023) or consistent face identities (Li et al., 2024; Wang et al., 2024b; Yan et al., 2023; Ruiz et al., 2024) across multiple generations. However, most existing approaches lack support for conditioning both characters and environments separately, often requiring predefined masks for generating characters (Chen et al., 2024; Lugmayr et al., 2022; Yang et al., 2023), with the environment remaining identical to the input image. This limitation makes it difficult to seamlessly integrate characters into different environments while ensuring both consistency and alignment with the input prompts. IP-Adapter (Ye et al., 2023) tackles this task by conditioning the generation on the environment and character images. However, IP-Adapter tends to over reconstruct the conditions, which causes interference between them. In this paper, we build our approach on IP-Adapter and propose an improved regional IP-Adapter with block drop, separating character and environment generation to enhance consistency.
@@
Large language models have demonstrated strong in-context learning capability (Brown, 2020), which enables them to solve diverse customized tasks based on human instructions and in-context examples. In the field of image generation, large language models have been employed for various tasks, such as image layout generator based on prompts (Cheng et al., 2024a; b), interactive multi-turn image generation (Zeqiang et al., 2023; Huang et al., 2024; Gong et al., 2023; Wang et al., 2023), and interleaved text and image generator (Team, 2024; Zhou et al., 2024a). Unlike these applications, our work focuses on distilling a specialized LLM, serving as a game engine, responsible for generating game mechanics, narratives and character interactions.
@@ 5 Results and Analysis
Quantitative Results We compare our regional IP-Adapter with block drop with previous approaches in maintaining environment consistency and character consistency. For all the approaches, we merge the character LoRA and LCM LoRA with the model to support fast inference and improve character consistency and provide an apples-to-apples comparison. As shown in Table 1, our approach consistently outperforms previous approach in maintaining environment consistency and character consistency, while achieving comparable performance in maintaining semantic alignment. Specifically, our approach significantly overtakes StoryDiffusion (Zhou et al., 2024b) by 0.047 in CLIP-IC, and 0.057 in DreamSimC for character consistency, and 0.035 in CLIP-IE, 0.065 in DINOE, and 0.058 in DreamSimE for environment consistency, demonstrating the effectiveness of our approach. Besides, our approach also achieves comparable performance in maintaining semantic alignment, suggesting strong text following capabilities.
@@ 6 Conclusion
Acknowledgements
We thank Shiran Zada, Peyman Millanfar and Shlomi Fruchter for the thoughtful feedback.

### v2

@@ Abstract
Figure 1: An example of Unbounded. We follow the life of Archibus, the user’s custom wizard character. The user can interact with the generative game using natural language, and Archibus’ hunger, energy and fun meters update accordingly. A spontaneous and unconstrained story unfolds while the user plays, and the character can explore new environments with a myriad of possible actions and unexpected interactions. The game runs in interactive speeds, refreshing every second.
@@ 1 Introduction
In his work “Finite and Infinite Games: A Vision of Life as Play and Possibility” (Carse, 1986), James P. Carse introduces a distinction between two types of games. Carse defines finite games as those “played for the purpose of winning,” with boundaries, fixed rules, and a definitive endpoint. In contrast, infinite games are “played for the purpose of continuing the play,” with no fixed boundaries and evolving rules. Traditional video games are, inherently, finite games due to the limitations of programming and computer graphics. For example, game mechanics have to be fully pre-defined in the programming language and graphics assets usually have to be pre-designed (modulo traditional procedural generation which still grapples with structural limitations). This allows for only a finite, and sometimes predefined, set of actions and paths that can be taken. They also feature predefined rules, boundaries, and win conditions, which run counter to the infinite game definition.
@@
Recent advances in generative models have been impressive. We hypothesize that these developments have finally opened up the possibility of creating the first generative infinite video game - an infinite game that is also fully subsumed in generative models with no logic or graphics being controlled by other, more traditional, processes. Two advancements have made this achievable: (1) large language models (LLMs) that can encode persistent video game mechanics (e.g. interactions with the game environment or characters, character state tracking, object permanence), generate interactive stories, and produce spontaneous, and sometimes emergent, behaviors; and (2) visual generative models capable of producing high-quality images that follow prompts. In this work, we present Unbounded, what we believe to be the first interactive generative infinite game, where all game behaviors and graphics are generated by AI models, transcending the constraints of hard-coded systems. Unbounded draws inspiration from sandbox life simulations and digital pet games such as Little Computer People, The Sims and Tamagotchi. It incorporates elements from tabletop roleplaying games like Dungeons & Dragons, which offer unconstrained storytelling experiences that have been unattainable in video games.
Some useful prior work has proposed ideas that are related to our concept of generative infinite games. Gaudl et al. (2018) proposes fluidic games that merge gameplay with game design, automated game design (Cook et al., 2016) explores automated systems that can help in making new games, and work on AI-based games (Treanor et al., 2015) proposes design techniques for games that put AI at the forefront of user experience. These concepts are different from our definition of generative infinite games since they emphasize structured exploration within predefined game parameters and tools, where the design space is extensive yet ultimately bounded. In contrast, infinite games, as we envision them, aim to continuously evolve beyond predefined structures, with generative models dynamically creating both content and mechanics in real-time, enabling an open-ended gameplay experience with no fixed endpoint or limitations on possible interactions. There is also important work on using machine learning models as components that generate parts of a game. Agarwal (2024) generates novel interactions, Sun et al. (2023) generates new mechanics and items as well as some graphical assets, and there is also work on orchestration of different models for game design (Liapis et al., 2019). These works do not generate all components of the game. In contrast, in our work, all of the game mechanics, characters, environments, narrative, and graphics are fully produced by generative models. This is more similar in spirit to recent work Genie (Bruce et al., 2024) and GameNGen (Valevski et al., 2024), although in contrast to these works that mainly generate platformers with diffuse mechanics or regenerate behaviors of one pre-existing game, our work proposes an open-ended narrative experience with stable game mechanics enabled by an LLM-based game engine.
@@ 2 Related Work
Controllable text-to-image generation becomes a key research direction in diffusion model applications, enabling diverse ways to guide the generation process. For instance, ControlNet (Zhang et al., 2023) introduces conditioning mechanisms that utilize control signals such as depth maps, poses, edges, and segmentation maps to guide image generation. Other works focus on layout control using bounding boxes to control object placement within the generated images (Li et al., 2023; Shin et al., 2022). Beyond these control signals, another major area of research involves personalization, where the goal is to generate consistent characters (Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Avrahami et al., 2024) or consistent face identities (Li et al., 2024b; Wang et al., 2024b; Yan et al., 2023; Ruiz et al., 2024) across multiple generations. However, most existing approaches lack support for conditioning both characters and environments separately, often requiring predefined masks for generating characters (Chen et al., 2024; Lugmayr et al., 2022; Yang et al., 2023), with the environment remaining identical to the input image. This limitation makes it difficult to seamlessly integrate characters into different environments while ensuring both consistency and alignment with the input prompts. IP-Adapter (Ye et al., 2023) tackles this task by conditioning the generation on the environment and character images. However, IP-Adapter tends to over reconstruct the conditions, which causes interference between them. In this paper, we build our approach on IP-Adapter and propose an improved regional IP-Adapter with block drop, separating character and environment generation to enhance consistency.
@@
Large language models have demonstrated strong in-context learning capability (Brown, 2020), which enables them to solve diverse customized tasks based on human instructions and in-context examples. In the field of image generation, large language models have been employed for various tasks, such as image layout generator based on prompts (Cheng et al., 2024a; b), interactive multi-turn image generation (Zeqiang et al., 2023; Huang et al., 2024; Gong et al., 2023; Wang et al., 2023b), and interleaved text and image generator (Team, 2024; Zhou et al., 2024a). Unlike these applications, our work focuses on distilling a specialized LLM, serving as a game engine, responsible for generating game mechanics, narratives and character interactions.
Game Generation.
In the field of Procedural Content Generation (PCG), diverse approaches have been explored to create dynamic game content (Summerville et al., 2018). Early works employed concept maps (Treanor et al., 2012), Markov Chains (Snodgrass & Ontañón, 2014), Bayes Nets (Guzdial & Riedl, 2016), and LSTMs (Summerville & Mateas, 2016) to simulate interactive game environments. Recent research has advanced to using Generative Adversarial Networks (GANs) for generating game levels or dynamic environments (Volz et al., 2018; Kumaran et al., 2019; Schubert et al., 2021; Kim et al., 2020), exploiting the use of diffusion models for game generation (Zhou et al., 2024b), utilizing LLMs to design and generate the game environments or mechanics (Sudhakaran et al., 2024; Todd et al., 2023; Nasir & Togelius, 2023; Hu et al., 2024; Zala et al., 2024; Anjum et al., 2024; Chung & Kreminski, 2024; Chung et al., 2024) and employing LLMs to generate items and narratives, as well as using diffusion models for scene generation in 1001 Nights (Sun et al., 2023). One major feature of this body of work is that AI is often employed to help in game design, or to generate one component of the game, for example environments or interactions. Our work aims to fully generate all behaviors of the game, including all the game’s graphics, characters, environments, and narrative, using generative models.
Bruce et al. (2024) explore synthesizing a completely new interactive game with a video diffusion model conditioned on previous game frames and user actions. This work achieves impressive initial results for this direction, at present it has limitations on the scope of the games being generated, with mostly 2D platformers being represented. This differs from our work since these are not infinite games. We also believe that generating compelling infinite games by only modeling pixels is a hard problem and hence we opt to use language models to generate all the open-ended game mechanics instead. Valevski et al. (2024) train a diffusion model to be a game engine running Doom, and achieve consistent generation of game mechanics by only modeling pixels. Nevertheless, this work currently fits the diffusion model to a single finite game that already exists.
Besides generating the game content, there has been exploration in letting AI take different roles (e.g., competitor, designer, or teammate) in games (Zhu et al., 2021; Gallotta et al., 2024; Pell, 1992; Agarwal et al., 2023), and a system with continuous creativity for automated game design for games that are played on a grid of up to 8x8 squares is built in  Cook (2022). Alongside generating visual game environments, another line of research builds narrative text games with conditional language generation models to complete a knowledge graph (Ammanabrolu et al., 2020) or utilizes LLMs or other AI tools to generate the game (Li et al., 2024a; Latitude Inc., 2023; Wang et al., 2023a) or act in different roles in the game (Zhu et al., 2023; Kreminski et al., 2020; Cui et al., 2023; Zhou et al., 2023; Dambekodi et al., 2020). Our work diverges from these by (1) tackling both generation of all game mechanics, graphics, characters, environments, and narrative of the game using generative models (2) introducing personalization into generative games, where a user can insert a custom character and personalize the game world and initial story arc, and (3) achieving interactive real-time refresh speed through our innovations in the vision and language domains.
@@ 5 Results and Analysis
Quantitative Results We compare our regional IP-Adapter with block drop with previous approaches in maintaining environment consistency and character consistency. For all the approaches, we merge the character LoRA and LCM LoRA with the model to support fast inference and improve character consistency and provide an apples-to-apples comparison. As shown in Table 1, our approach consistently outperforms previous approach in maintaining environment consistency and character consistency, while achieving comparable performance in maintaining semantic alignment. Specifically, our approach significantly overtakes StoryDiffusion (Zhou et al., 2024c) by 0.047 in CLIP-IC, and 0.057 in DreamSimC for character consistency, and 0.035 in CLIP-IE, 0.065 in DINOE, and 0.058 in DreamSimE for environment consistency, demonstrating the effectiveness of our approach. Besides, our approach also achieves comparable performance in maintaining semantic alignment, suggesting strong text following capabilities.
@@ 6 Conclusion
Acknowledgments
We thank Shiran Zada, Peyman Millanfar, Shlomi Fruchter, Michael Goin and Matthew Guzdial for the thoughtful feedback and discussion.